# Comparative analysis of diverse methodologies for portfolio optimization leveraging quantum annealing techniques

### v2

@@ II Portfolio optimization formulation
the given formulation in Markowitz’s mean-variance model is a binary constrained mean-variance model. This model addresses the binary portfolio optimization problem, where the variables are limited to values of 0 or 1, $\mu_{i}$ is the expected return of assets, $\sigma_{ij}$ is the risk of assets, and $B$ is the full budget.
The formulation of a particular portfolio optimization depends on the given market conditions. To adapt a portfolio optimization problem for quantum annealing, it is necessary to transform the formulation into a Quadratic Unconstrained Binary Optimization (QUBO) problem through various conversions, such as binary, unary, or sequential encoding. [89, 65]. The problem is encoded into a problem Hamiltonian $H_{P}$, which is defined by pauli matrices
@@ III.3 RF-QA
Algorithm
Scaling exponent
Standard QA
0.70
Ferro Coupler QA
0.75
Mixed Coupler QA
0.70
Anti-ferro Coupler QA
1.10
Inhomogeneous QA
0.85
$RFQA_{M}$
1.17
$RFQA_{D}$
1.16
CD-QA
0.89
@@ IV Reverse annealing
Algorithm
Scaling exponent
Standard QA
1.00
Reverse Standard QA
0.34
Reverse Ferro Coupler QA
0.35
Reverse Mixed Coupler QA
0.34
Reverse Anti-ferro Coupler QA
0.35
Reverse $RFQA_{M}$
0.09
Reverse $RFQA_{D}$
0.09

### v3

@@ II Portfolio optimization formulation
the given formulation is a constrained mean-variance model, where $\mu_{i}$ is the expected return of assets, $\sigma_{ij}$ is the risk of assets, and $B$ is the full budget.
The formulation of a particular portfolio optimization depends on the given market conditions. To adapt a portfolio optimization problem for quantum annealing, it is necessary to transform the formulation into a Quadratic Unconstrained Binary Optimization (QUBO) problem, where the variables are limited to values of 0 or 1. The transformation can be made through various conversions, such as binary, unary, or sequential encoding. [89, 65]. The problem is encoded into a problem Hamiltonian $H_{P}$, which is defined by pauli matrices
@@ III.3 RF-QA
Algorithm Scaling exponent
Standard QA 0.70
Ferro Coupler QA 0.75
Mixed Coupler QA 0.70
Anti-ferro Coupler QA 1.10
Inhomogeneous QA 0.85
$RFQA_{M}$ 1.17
$RFQA_{D}$ 1.16
CD-QA 0.89
@@ IV Reverse annealing
Algorithm Scaling exponent
Standard QA 1.00
Reverse Standard QA 0.34
Reverse Ferro Coupler QA 0.35
Reverse Mixed Coupler QA 0.34
Reverse Anti-ferro Coupler QA 0.35
Reverse $RFQA_{M}$ 0.09
Reverse $RFQA_{D}$ 0.09